OpenStudy (anonymous):

Can I have a detailed explanation, so I can do this on my own please? I will fan and medal!! What is the exact area of a circle with a diameter of 19m?

OpenStudy (anonymous):

we have given that diameter = 19 m we know that radius = diameter/2 so that we get radius = 19/2 formula of the area of circle = pi*r^2 plug r = 19/2 in the above formula so the area of circle = pi*(19/2)^2 =pi(361/4) = 3.14*361/4 the area of circle =283.385 square meter.

OpenStudy (anonymous):

@KESHAVRP Ok, Thank you~!

OpenStudy (anonymous):

welcome.